OpenStudy (anonymous):

given f(x)= the square root of (1+3e^x), find f'(x) thanks :)

OpenStudy (zarkon):

\[f(x)=(1+3e^x)^{1/2}\] \[f'(x)=\frac{1}{2}(1+3e^x)^{-1/2}3e^x\]